OpenStudy (anonymous):

what is the exact value of cos(-5*pi over3)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf cos~\frac{-5π}{3} }\) \(\LARGE\color{blue}{ \bf cos~\frac{-5(180)}{3} }\) \(\LARGE\color{blue}{ \bf cos~(-300) }\) \(\LARGE\color{blue}{ \bf cos~(300) }\) \(\LARGE\color{blue}{ \bf cos~(360-60) }\) can you go on from here, or still need help ?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf cos~(A-B)=cosAcosB+sinAsinB }\)

OpenStudy (anonymous):

\[\cos \left( -\theta \right) =\cos \theta \] \[\cos \left( \frac{ 5 \pi }{ 3} \right)=\cos \left( 2 \pi-\frac{ \pi }{ 3 } \right)=\cos \frac{ \pi }{ 3 }=?\]